Please help me to solve this problem: 12-4k=-18+11k
Can you show me the steps to help me know how to do the problem.

Answer by stanbon(75887)

12-4k=-18+11k
Add 4k to both sides to get:
12 = -18 + 15k
Add 18 to both sides to get:
30 = 15k
Divide both sides by 15 to get:
k = 2
------------
Cheers,
Stan H.